### v2

# Can Language Models Recognize Convincing Arguments?

Abstract
The capabilities of large language models (LLMs) have raised concerns about their potential to create and propagate convincing narratives. Here, we study their performance in detecting convincing arguments to gain insights into LLMs’ persuasive capabilities without directly engaging in experimentation with humans. We extend a dataset by Durmus and Cardie (2018) with debates, votes, and user traits and propose tasks measuring LLMs’ ability to (1) distinguish between strong and weak arguments, (2) predict stances based on beliefs and demographic characteristics, and (3) determine the appeal of an argument to an individual based on their traits. We show that LLMs perform on par with humans in these tasks and that combining predictions from different LLMs yields significant performance gains, surpassing human performance. The data and code released with this paper contribute to the crucial effort of continuously evaluating and monitoring LLMs’ capabilities and potential impact. (https://go.epfl.ch/persuasion-llm)
Can Language Models Recognize Convincing Arguments?
Paula Dolores Rescala1, Manoel Horta Ribeiro1, Tiancheng Hu2, Robert West1 1 EPFL, 2 University of Cambridge Correspondence: manoel.hortaribeiro@epfl.ch
1 Introduction
As LLMs rise in capacity and popularity, so has the concern that they may help create and propagate tailor-made, convincing narratives De Angelis et al. (2023); Buchanan et al. (2021). While “tailor-made misinformation” predates LLMs (DiResta et al., 2019), frontier models such as GPT-4, Claude 3, and Gemini 1.5 could add to the problem by allowing malicious actors to easily create diverse, personalized content (Bommasani et al., 2021; Goldstein et al., 2023) or enable the detection (and amplification) of existing content that would be particularly persuasive to individuals with specific demographics or beliefs (Broniatowski et al., 2018).
Previous work has found LLMs to be persuasive in the generative setting Simchon et al. (2024); Hackenburg and Margetts (2024); Breum et al. (2024); for example, Salvi et al. (2024) found that, when provided with personal attributes, GPT-4 outperformed crowdworkers in a debate setting. Yet, assessing models’ capacity to generate arguments requires continuous human experimentation as LLMs evolve, which can be time-consuming and resource-intensive. On the contrary, measuring a model’s capacity to detect content persuasive to specific demographics can be done quickly and without interaction with human subjects, making it a more efficient approach for benchmarking the persuasive capabilities of LLMs.
Figure 1: Our approach to study LLMs’ persuasiveness capabilities. We measure to which extent LLMs can reproduce human judgments on the quality and persuasiveness of arguments. Suppose LLMs can predict users’ positions on stances (e.g., The death penalty should be legal) before and after reading a debate and judge who the better debater was. In that case, they would be well suited to power personalized misinformation and propaganda.
Present Work.
[…]
  • RQ1: judge the quality of arguments and identify convincing arguments and humans?
  • RQ2: judge how demographics and beliefs influence people’s stances on specific topics?
  • RQ3: determine how arguments appeal to individuals depending on their demographics?
To investigate these questions, we extend a dataset collected by Durmus and Cardie (2018) from a defunct debate platform (debate.org). We annotate 833 politics-related debates with clear propositions, such as “The electoral college should remain unchanged.” Each debate contains arguments for (“Pro”) and against (“Con”) the proposition, along with votes from debate.org participants indicating the winning side. Importantly, the dataset includes demographic information of the voters as well as their stances on 48 so-called “big issues”. For 121 debates with 751 votes on three of the most prominent topics in the dataset, we obtained crowdsourced labels to compare the capabilities of LLMs to those of humans. Then, using this enriched dataset, we evaluate the performance of four LLMs (GPT-3.5, GPT-4, Llama 2, and Mistral 7B) on three tasks: 1) identifying the side with more convincing arguments (RQ1); 2) predicting individuals’ stances on specific propositions before the debate, given their demographic and basic belief information (RQ2); and 3) predicting individuals’ stances on specific propositions after the debate, given their demographic and basic belief information (RQ3). Figure 1 illustrates our approach.
Our key finding is that LLM exhibits human-like performance across the three proposed tasks. In judging the better debater (RQ1), GPT-4 (Accuracy: 60.50%) is as good as an individual voter in the dataset (Accuracy: 60.69%;). When predicting users’ stances on specific issues before and after reading the debate (RQ2 and RQ3), LLMs again perform similarly to humans. For instance, in the before the debate scenario (RQ2), Mistral yields an accuracy of 42.27%, whereas crowdworkers achieve 39.86% (random guessing would yield 33.3% accuracy). However, zero-shot prediction with LLMs still underperforms a supervised machine learning model [XGBoost (Chen and Guestrin, 2016)], which achieves 58.25% accuracy in cross-validation. Nevertheless, stacking the predictions of LLMs and using them as features in a supervised learning setting reduces the performance gap (45.91%).
Overall, our work contributes to the growing body of research on the societal impact of LLMs (Bommasani et al., 2021; Solaiman et al., 2023; Weidinger et al., 2023, inter alia). We shed light on the potential misuse of LLMs by investigating their ability to detect persuasive content tailored to specific demographics.
2 Related Work
We review related work in three broad directions closely related to the tasks proposed.
Demographics, beliefs, and persuasion.
Demographics have long been known to impact people’s political beliefs and attitudes. Group-level demographic factors such as race, religion, and education shape individuals’ perspectives on various political issues and voting behavior in the U.S. context (Campbell et al., 1960; Erikson and Tedin, 2019). For example, 78% of Black, 72% of Asian, and 65% of Hispanic workers see efforts on increasing diversity, equity, and inclusion at work positively, compared to 47% of White workers (Minkin, 2023). Similarly, previous work indicates that persuasion depends on the message recipients’ existing values and that individual differences can influence persuasion (O’Keefe, 2015). For instance, Hirsh et al. (2012) demonstrated that tailoring messages to different personality traits can make them more persuasive; Orji et al. (2015) showed that men and women differ significantly in their responsiveness to the different persuasive strategies. Closer to the work at hand, Durmus and Cardie (2018) and Al Khatib et al. (2020) showed considering demographic characteristics can enhance the prediction of argument persuasiveness. However, the extent to which LLMs can utilize demographic characteristics in persuasiveness judgment remains underexplored. In this work, we examine how LLMs can capture the correlations between demographics and beliefs (RQ2) and how personal attributes determine the persuasiveness of arguments (RQ3).
Argument Mining and Argument quality.
Defining argument quality is no easy task, or as persuasion scholars O’Keefe and Jackson (1995) have put it: “there is no clear general abstract characterization of what constitutes argument quality.” An argument may be deemed good due to its effectiveness in convincing people (O’Keefe and Jackson, 1995), its cogency from individually accepted premises that lead to a conclusion (Johnson and Blair, 2006), or its reasonableness in contributing to resolving a disagreement (Walton, 2005). Over recent decades, there has been significant interest in automatically extracting arguments from text (Habernal and Gurevych, 2016, 2017; Swanson et al., 2015, inter alia), as detailed in surveys like (Cabrio and Villata, 2018; Lawrence and Reed, 2020). Additionally, research has explored computational argument quality and persuasiveness analysis (Habernal and Gurevych, 2016; Tan et al., 2016; Wachsmuth et al., 2017, inter alia). Contemporary works begin to explore the potential of LLMs in argument quality judgement Mirzakhmedova et al. (2024); Wachsmuth et al. (2024). Our work complements existing work by examining the extent to which LLMs can identify higher-quality arguments in a debate setting (RQ1) and determine argument effectiveness across individuals with different demographics and beliefs (RQ3).
Personalized misinformation and propaganda.
Microtargeting or “personalized persuasion” refers to tailoring the language or content of messages to individuals based on their characteristics (e.g., demographics and prior political beliefs) to make them maximally persuasive. Evidence on the effect of microtargeting is mixed (Guess and Coppock, 2020; Coppock et al., 2020; Matz et al., 2017; Tappin et al., 2023), which has led Teeny et al. (2021) to propose that research on microtargeting should move from “does microtargeting work?”, to “when micro-targeting works?” At the same time, the increasing popularity and capabilities of LLMs have raised concerns that they may not only make microtargeting cheaper and more effective but also enable new forms of “microtargeting” misinformation and propaganda, such as through personalized chatbots—see Goldstein et al. (2023) for a comprehensive discussion. These concerns are corroborated by recent studies suggesting that LLMs are capable of generating messages perceived as equally or more persuasive than humans (Bai et al., 2023; Durmus et al., 2024); that they can personalize messages to make them more persuasive (Simchon et al., 2024); and that LLMs can successfully persuade humans in debates by exploiting their personal traits Salvi et al. (2024). One key drawback, though, is that these studies typically involve large and expensive experiments that cannot easily be replicated when a new LLM is released or explore the large hyperparameter space of existing models (e.g., prompting strategy and decoding algorithm). In our work, we argue that we could instead evaluate the effectiveness of LLM in determining whether someone of a specific set of demographic characteristics would find an argument convincing (RQ3) and view this as a proxy of the LLM’s ability to perform political microtargeting.
Social sensing.
Prediction tasks where individuals are asked to determine the preferences and opinions of others have been broadly referred to as social sensing Galesic et al. (2021). Previous work using this approach has shown that human social sensing outperformed traditional polling in forecasting elections Galesic et al. (2018) and that the approach is useful in predicting disease outbreaks Christakis and Fowler (2010). Here, the tasks associated with RQ2 and RQ3 are, in their essence, social sensing tasks, as we ask LLMs (and crowdworkers) about the preferences and opinions of others. Although informative, predictions obtained through human social sensing are known to be subjected to biases Ross et al. (1977); Chambers and Windschitl (2004), and therefore, it is possible that so are predictions obtained through LLM social sensing.
Original Title Hand-written Proposition
A Debate On The Electoral College The electoral college should remain unchanged.
Gay Marriage Should Be Legal Gay marriage and equal rights.
US Hegemony U.S. hegemony is desirable.
Abortion Abortion should be illegal.
Table 1: Examples of titles used for debates in the dataset and the corresponding manually written propositions we created to replace them.
3 Data
Data for this study was collected by Durmus and Cardie (2018) from an online debate platform (debate.org; no longer operational). The platform allowed users to participate in and vote on debates covering a breadth of topics, including politics, religion, and science. Each debate within the dataset consists of multiple rounds, each round with an argument from both the “Pro” and “Con” perspectives. Users on the platform could vote on various aspects of the debate, such as which side they believed provided a more convincing argument. The raw dataset contains 78,376 debates, 45,348 users, and 195,724 votes. Each user has corresponding demographic information, such as gender and age, as well as their stances on 48 so-called “big issues,” such as abortion, capital punishment, and national health care (see Appendix B for details). Nevertheless, most demographic data is missing from the dataset. Most important to the work at hand, voters had to indicate which side: 1) Made more convincing arguments; 2) They agreed with before the debate; 3) They agreed with after the debate. We measure LLMs’ capacity to recognize convincing arguments by predicting the responses to these three questions, each of which could be answered “Pro,” “Con,” or “Tie.” Note that predicting question #1–#3 corresponds to our research questions RQ1RQ3.
Although each debate in the dataset has a corresponding title indicative of its content, these titles are user-defined and do not always take the form of a proposition. As a result, it is not always clear from reading the debate title alone what the “Pro” and “Con” stances are. Hence, we contribute clear, manually written propositions for 833 debates that (1) were categorized under “Politics,” (2) contained at least 300 total tokens (tokens are counted using the tiktoken library with the GPT-3.5-turbo model encoding), (3) contained at least two complete rounds, (4) The debater who spoke the most in the debate did not speak more than 25% more than the other debater, (5) the debate had at least three votes. We discarded an additional 199 debates that fulfilled the aforementioned criteria but were troll debates (e.g., just profanity toward the other debater), incorrectly categorized as Politics, or impossible to paraphrase into a proposition (see Table 1 for examples).
PoliProp [PP].
We study these 833 annotated debates, considering all votes ($n=4$,$871$) in these debates for users with no more than five missing values in demographic information ($4$,$871$ out of $7{,}797$). We also trimmed each debate in the dataset larger than the smallest context window (4096 tokens) among LLMs considered. Trimming is done by removing one round at a time from the end of the debate until the token count is small enough, an approach that equally penalizes both debaters (unlike simply removing tokens at the end of the debate). Hereafter, we call this the PoliProp dataset.
PoliIssues [IS].
We also separately consider all debates on abortion ($n=50$), gay marriage ($n=51$), and capital punishment ($n=31$), the most prominent topics in the dataset. Given that debates within the three themes are similar, we use this data to compare LLM performances with traditional machine learning methods, predicting participants’ votes using their demographic and stances on big issues as features. To obtain a human baseline, we collect crowdsourced labels using Amazon Mechanical Turk (MTurk) for each of the 751 votes cast on these 121 debates. Crowdworkers are essentially presented with the same questions as the LLMs. Given a debate, we ask who gave the better arguments. Given a set of characteristics by a voter as well as the debate, we ask whether the voter would have agreed with the proposition before and after reading the debate. Hereafter, we call this dataset the PoliIssues dataset. For more information on crowdsourcing, see Appendix C.
Figure 2: Prompt structure used in RQ1.
Figure 3: Prompt structure used in RQ2.
Figure 4: Prompt structure used in RQ3.
4 Methods
LLMs considered.
For this study, we compare the performance of two open-source LLMs, namely Mistral 7B (Mistral-7b-Instruct-v0.1) and Meta’s Llama 2 70B (Llama-2-70b-chat), with OpenAI’s closed-source GPT-3.5 (gpt-3.5-turbo-1106) and GPT-4 (gpt-4-0613). We use the standard temperatures for each model.
Prompting.
We follow Staab et al. (2023) to develop our prompt: each had a system role, context, question, and constraint. We experimented with different structures and found that, overall, the structure mattered little as long as the wording was clear and concise. Since we had three research questions to answer, we had three prompt structures that combined the debate proposition, the debate itself, and user demographics. We show the prompt structure used for RQ1 in Figure 2. All prompts indicated that LLMs should respond only with the labels “Pro,” “Con,” or “Tie”. Nevertheless, many of the models failed to adhere to this instruction, necessitating post-processing to extract the actual answer. Generally, instances of incorrect responses involved the answer accompanied by additional spaces or punctuation or presented in a complete sentence format, such as: “Based on the given demographics, the person is most likely to agree with the ‘Con’ side in the debate.” We used heuristics to extract the responses in these cases. Additionally, there were occasions when the LLMs failed to produce any answer, resulting in responses akin to: “I cannot determine the person’s position in the debate without additional information.” We depict the remaining prompts in Figures 3 and 4 and provide further details in Appendix A.
Evaluation.
We evaluate the accuracy of language models by comparing the answers they provide with the ground truth data from PoliProp and PoliIssues. We obtain confidence intervals through bootstrapping. Besides considering each LLM individually, we also consider the performance of stacked LLM predictions, obtained by using the output of different LLMs as features in a supervised machine learning model Hastie et al. (2009).
Baselines/Benchmarks
We interpret the LLM accuracies by establishing the following baselines and benchmarks as metrics of comparison:
  • Random; (RQ1—RQ3) Since there are three possible stances (Pro, Con, and Tie) for any given task and each debate and voter pair, the random baseline has an accuracy of 33.3%.
  • Majority; (RQ1) For RQ1, the ground truth was established by aggregating the votes for who made more convincing arguments in each debate through a simple majority vote. The Majority benchmark is the percentage of users in our dataset that agreed with the computed ground truth for this question.
  • MTurk; (RQ2–RQ3) We crowdsourced the tasks for each research question for the PoliIssues dataset, obtaining a human equivalent answer to the questions we asked the LLMs. These are detailed in Appendix C.
  • XGBoost; (RQ2) For each issue (abortion, gay marriage, capital punishment) in the PoliIssues dataset, we train a Gradient Boosting classifier to predict the stance of a user as in RQ2. We train one model separately per issue since labels are not equivalent (e.g., Pro-abortion differs from Pro-capital punishment), but we report the aggregated accuracy.
# Question Model Dataset Accuracy (%) 95% Confidence Interval
1 RQ1 Llama 2 PP 24.91 (20.65, 26.53)
2 Mistral 7B PP 37.69 (32.89, 39.26)
3 GPT-3.5 PP 42.74 (39.38, 46.1)
4 GPT-4 PP 60.50 (57.26, 63.87)
5 Stacked PP 61.94 (58.54, 65.34)
6 Majority PP 60.69 (59.56, 61.79)
7 Random PP 33.33
8 RQ2 Llama 2 IS 41.56 (38.16, 45.1)
9 Mistral 7B IS 41.39 (38.4, 44.86)
10 GPT-3.5 IS 41.73 (38.52, 44.98)
11 GPT-4 IS 42.82 (39.59, 46.53)
12 MTurk IS 39.32 (35.89, 42.88)
13 Stacked IS 45.91 (40.02, 51.81)
14 XGBoost IS 55.34 (50.45, 60.22)
15 Random IS 33.33
16 RQ3 Llama 2 IS 41.24 (37.2, 44.02)
17 Mistral 7B IS 42.28 (32.06, 38.28)
18 GPT-3.5 IS 38.97 (35.41, 42.11)
19 GPT-4 IS 44.38 (40.91, 47.73)
20 MTurk IS 39.86 (36.44, 43.42)
21 Stacked IS 46.86 (41.17, 52.55)
22 Random IS 33.33
Table 2: Key results for RQ1–RQ3. We show that LLMs perform on par with humans across various tasks related to recognizing convincing arguments. When stacked using a logistic regression, LLMs outperform humans in predicting stances on prepositions before and after the debate (RQ2, RQ3). The random baseline has accuracy of 33.33% for all settings.
5 Results
Judging argument quality (RQ1).
Considering the PoliProp dataset [PP], we summarize the accuracy of the different LLMs and baseline methods in determining argument quality in Table 2 (rows #1–#7). We find a substantial performance gap between GPT-4 (60.50% accuracy) and the other models, e.g., Llama 2, which performs worse than random guessing (24.91%). GPT-4 performance is similar to human performance, as measured by the agreement of any individual vote with the remaining votes in each debate (Majority; 60.69%).
Correlating beliefs and demographic characteristics with stances (RQ2).
Considering the PoliIssues dataset [IS], we summarize the accuracy of the different LLMs and baseline methods in correlating beliefs and demographic characteristics with stances on Table 2 (rows #8–#15). Here, the accuracy range of different LLMs is much more narrow, ranging from 41.39% (Mistral) to 42.82% (GPT-4). Most important, however, is that the performance of LLMs is similar to that of crowdworkers (39.32%; MTurk).
Recognizing convincing arguments (RQ3).
Again, considering the PoliIssues dataset [IS], we summarize the accuracy of the different methods in recognizing users’ opinion after reading the debate on Table 2 (rows #16–#22). Different models perform similarly on the task and similar to crowdworkers, e.g., GPT-4: 44.38% of vs. crowdworkers: 39.86%.
LLMs vs. supervised learning.
Considering RQ2, we train a Gradient Boosting classifier to predict stances given user traits (row #14). We run a 20-fold cross-validation and report the mean accuracy. This model performs significantly better than LLMs at predicting stances (Accuracy: 58.25%; 95% CI: [54.02, 62.47]).
Model Big Issues Reasoning Accuracy (%) 95% CI
Llama 2 False False 41.30 (37.92, 44.38)
False True 40.05 (36.48, 43.3)
True False 38.92 (34.81, 41.51)
True True 37.38 (29.07, 35.17)
Mistral 7B False False 40.67 (37.2, 44.02)
False True 41.83 (38.04, 44.98)
True False 40.60 (36.96, 44.14)
True True 40.91 (36.6, 43.18)
GPT-3.5 False False 42.94 (39.83, 46.17)
False True 39.45 (36.0, 42.58)
True False 41.80 (38.28, 45.1)
True True 37.80 (34.57, 41.03)
GPT-4 False False 42.70 (39.47, 46.17)
False True 43.30 (39.95, 46.65)
True False 42.46 (39.11, 45.93)
True True 45.03 (41.51, 48.09)
Table 3: We repeat the analysis to answer RQ2 using the PoliIssues dataset but varying the prompt, either by considering big issues in the prompt (Big Issues) or by asking the LLM to reason before answering the question (Reasoning). The scenario without ‘Big issues’ or ‘Reasoning’ corresponds to lines #8–10 in Table 2,
Sensitivity to prompt.
We study whether the results obtained were sensitive to the prompt used by re-running the analysis from RQ2 on the PoliIssues dataset. For each model, we rerun the analysis considering the “big issues” in user profiles and/or asking for models to reason before answering. [^1] The prompt constraint was changed to Evaluate step-by-step the data given in the proposition before coming to an answer. Provide your reasoning for selecting an answer and then give your answer in the form of ‘Pro,’ ‘Con,’ or ‘Tie’ without using other words or punctuation. Provide your response in the following format: ‘Reasoning: your reasoning goes here. Answer: your answer goes here.’ Results are shown in Table 3. Overall, we find that the results are not sensitive to the experimented changes.
Figure 5: Inter-annotator agreement for different models in RQ2.
Stacking LLMs.
While the performance of language models is similar in RQ2, we find that their inter-annotator agreement is quite low (Cohen’s $\kappa$ is smaller than 0.2 for most pairs of models, see Figure 5). This is surprising since, upon our inspection of the reasoning different LLMSs’ provided for their answers, all made similar assumptions. Nevertheless, each model seems to perform well on a different subset of the debates. This motivated us to experiment with stacking LLMs, i.e., using the outputs of the different large language models outlined above as input to a simple logistic regression model. We find this strategy yields a small boost in accuracy in RQ1 (see row #5; Table 2), but a substantial one in RQ2 and RQ3 (see rows #13 and #21). Indeed, in this scenario, the accuracy is significantly better than crowdworkers for both research questions ($p<0.05$). Note that the accuracy reported for the stacked model is the average of a 20-fold cross-validation.
6 Discussion and Conclusion
Here we studied LLM’s persuasive capabilities by considering its ability to identify convincing arguments in general and for people with specific arguments. We argue that if LLMs can detect content that is highly persuasive to specific demographics, they may be used to detect and amplify tailor-made misinformation and propaganda. Our findings indicate that LLMs demonstrate human-level performance in (1) judging argument quality, (2) predicting users’ stances on specific topics given users’ demographics and basic beliefs, and (3) detecting arguments that would be persuasive to individuals with specific demographics or beliefs.
However, the overall human performance is not high in each of the three tasks [around 60% for (1), and around 40% for (2) and (3)], which could be due to the inherent difficulty of the tasks, as well as variance and randomness in the data. This does not necessarily imply that LLMs do not pose any additional risk of tailor-made misinformation in the future. It is plausible that with access to more personal information about an individual, such as personality traits, LLMs could perform better at detecting persuasive arguments (Hirsh et al., 2012). Nevertheless, it is important to consider that the more fine-grained the target, the harder and more costly it becomes to reach the targeted population, and the cost-benefit analysis is not straightforward (Tappin et al., 2023).
One hypothesis that could explain the relatively low accuracy for both LLMs and human performance is that these demographic questions and big-issue stances may not be highly relevant for the task, as suggested by Hu and Collier (2024). However, this is contradicted by the fact that a supervised XGBoost model trained with these factors yields much better results. Interestingly, stacking various LLM predictions yields performance closer to XGBoost. This indicates that while an individual LLM may not excel at detecting persuasive arguments for an individual, combining the predictions of several LLMs could achieve much more competitive performance (perhaps because each LLM’s biases differ). Consequently, LLMs can potentially detect highly effective tailored misinformation and propaganda, particularly in a multi-agent setting Schoenegger et al. (2024).
Limitations
Our dataset, from debate.org, may not be representative of the general population. The demographics of individuals opting to participate in online debates are likely skewed compared to the U.S. population and even more so globally. Additionally, we could not test the language models on non-English data due to data access limitations. However, recent research has shown that language models’ performance is considerably lower for non-English languages, especially low-resource ones (Ahuja et al., 2023). Consequently, it is plausible that the risk of misuse for microtargeting in non-English settings is currently lower. Nevertheless, as language models continue to improve, it is crucial to expand this line of research to a wider range of languages and demographics to ensure a comprehensive understanding of the risks associated with personalized persuasion. It is also essential to conduct empirical studies to understand whether LLMs are, in fact, being used for persuasion in online settings (e.g., in social media platforms). Another limitation of the work at hand is that the LLMs studied might have seen content from debate.org in their training data. To address this concern, we queried 100 debate excerpts from the dataset using GPT4 and couldn’t obtain complete samples. Yet, this is not sufficient to rule out this possibility.
Ethical Considerations
In this study, we employ demographic and belief-related questions drawn from datasets that are publicly accessible and have been anonymized before release. It is crucial to emphasize the importance of responsible development and deployment of LLMs and the need for ongoing research into mitigating their potential risks Bommasani et al. (2021). Our work can inform the development of safeguards and countermeasures against the misuse of LLMs for personalized misinformation and propaganda.
In this study, we employed crowdsourcing to evaluate the persuasiveness of debates. We paid crowd workers, all based in the U.S., at a rate of \$12.00 per hour, higher than the federal minimum wage in the United States.
Appendix A Prompts
In all tasks conducted for this study, the LLMs were prompted to respond to questions using only one of the options: "Pro," "Con," or "Tie," without any additional words or punctuation. Nevertheless, many of the models failed to adhere to this instruction, necessitating post-processing to extract the actual answer. Generally, incorrect responses involved the answer being accompanied by additional spaces or punctuation or presented in a complete sentence format, such as: “Based on the given demographics, the person is most likely to agree with the ‘Con’ side in the debate.” [^2] There were occasions when the LLMs failed to produce any answer, resulting in responses akin to: “I cannot determine the person’s position in the debate without additional information.” To extract the answer in these cases, we used a simple Regex expression, finding the first occurrence of the words “Pro,” “Con,” or “Tie.”
Table 4 shows what percentage of responses followed instructions and the corresponding percentage from which answers could be successfully extracted for the PoliProp dataset across each question. The answer extracted percentage indicates the highest achievable accuracy for each model in its results. Notably, both open-source models encountered challenges in complying with the instructions, with particular difficulty in addressing Q3.
Question Model Correct Form (%) Answer Extracted (%)
1 GPT-3.5 99.88 100.00
1 GPT-4 99.06 100.00
1 Llama 2 0.00 95.08
1 Mistral 62.76 95.55
2 GPT-3.5 99.71 99.77
2 GPT-4 99.80 99.80
2 Llama 2 0.00 97.17
2 Mistral 67.14 100.00
3 GPT-3.5 99.82 99.94
3 GPT-4 99.61 99.98
3 Llama 2 0.04 97.17
3 Mistral 17.19 81.48
Table 4: Some models had difficulty following instructions and giving the answer in the correct form of either "Pro," "Con," or "Tie." In this table, we see what percentage of the answers were given in the correct form and what percentages contained an answer after processing the result for the PoliProp dataset.
Appendix B Demographics and Big Issues
The dataset by Durmus and Cardie (2018) contained the following demographic information about participants: birthday, education, ethnicity, gender, income, party, political ideology, religious ideology.
It also contained participants’ opinions on so-called “big issues.” They were: abortion, affirmative action, animal rights, Barack Obama, border fence, capitalism, civil unions, death penalty, drug legalization, electoral college, environmental protection, estate tax, European Union, euthanasia, federal reserve, flat tax, free trade, gay marriage, global warming exists, globalization, gold standard, gun rights, homeschooling, Internet censorship, Iran-Iraq war, labor union, legalized prostitution, Medicaid and medicare, medical marijuana, military intervention, minimum wage, national health care, national retail sales tax, occupy movement, progressive tax, racial profiling, redistribution, smoking ban, social programs, social security, socialism, stimulus spending, term limits, torture, United Nations, war in Afghanistan, war on terror, and welfare.
Appendix C Crowdsourcing
We recruited participants for our study through Amazon Mechanical Turk between December 2023 and March 2024, requiring that they be 18+ years old, located in the US, and have a master’s qualification provided by Amazon. The study was paid \$2.25 and had a median completion time of 11 minutes, corresponding to a pay rate of about \$12.00/hour. To ensure the quality of answers, we asked users to justify their responses to each question. We then manually assessed the responses and considered them to be high-quality. We reproduce the crowdsourcing questions on the next page. We also provide an example justification below.
  • S#1: Being a Democrat and to a lesser extent white and female all correlate with being pro-LGBTQ.
  • S#3: The con side goes off on an unusual, libertarian leaning bend that probably just wouldn’t appeal to this type of person who would simply connect with the pro side more.
  • S#3: The con side argues less directly about this particular topic and more about some kind of libertarian; the state should have nothing to do with any of this kind of thing, which just isn’t as compelling as the pro side making clear why gay people should be integrated into the current system. The con side also repeatedly appeals to some really weak slippery slope stuff and doesn’t engage well with how the pro side responds.
Subtask 1
Read the following proposition, i.e., a statement that affirms or denies something.
Proposition: Gay marriage should be legal.
Consider an individual with the following demographic characteristics.
  1. 1.
    Education: Graduate Degree
  2. 2.
    Gender: Female
  3. 3.
    Party: Undecided
  4. 4.
    Political Ideology: Progressive
  5. 5.
    Religious Ideology: Christian
  • In your opinion, would this person agree (Pro), disagree (Con), or be neutral or undecided (Tie) with the proposition?
  • Write a brief justification for your answer. A sensible justification is required for your HIT to get approved.
Subtask 2
Consider the following debate on the proposition, where one individual argues for the proposition (Pro) and another against (Con).
Proposition: Gay marriage should be legal.
[debate]
Consider an individual with the following demographic characteristics.
  1. 1.
    Education: Graduate Degree
  2. 2.
    Gender: Female
  3. 3.
    Party: Undecided
  4. 4.
    Political Ideology: Progressive
  5. 5.
    Religious Ideology: Christian
  • Given this information, what stance do you think this person would take on the above proposition after reading the debate? Answer the same as before if you believe the debate had no effect on their opinion, and choose a different answer if you believe the debate had an effect on their opinion.
  • Write a brief justification for your answer. A sensible justification is required for your HIT to get approved.
Subtask 3
Again, consider the same debate on the proposition.
Proposition: Gay marriage should be legal.
[debate]
  • Disregarding your own point of view on the debate, please determine which debater you believe had more convincing arguments. The individual arguing for the proposition (Pro) or against it (Con)? If both were similarly convincing, indicate that it was a "Tie."
  • Write a brief justification for your answer. A sensible justification is required for your HIT to get approved.